# DiscoNeRF: Class-Agnostic Object Field for 3D Object Discovery

### v2

@@ 4.2 Ablation study
To validate the impact of the different components of our method, we perform an ablation study. We remove the matching $\gamma$ and instead match masks and objects by choosing the slot that maximizes the affinity $\alpha$ independently from each other, such that the same object slot can be assigned to several masks. This leads to a sharp drop in our evaluation metrics from $79.23\%$ to $57.04\%$ IoU, as the same object slot often captures several objects in Fig. 5(b).
We also ablate the regularization loss $\mathcal{L}_{TV}$ and compare the produced segmentations in Fig. 5(c). We observe a drop in IoU down to $75.53\%$ that can be attributed to the noisy object field learned without our additional regularization. Interestingly, we also observed in our experiments that the addition of $\mathcal{L}_{TV}$ helped prevent different objects from using the same object slot. We view this as a consequence of our regularization, which prevents the optimization from getting trapped in a local minimum. We further observe that without $\mathcal{L}_{TV}$, the predicted confidence remains low and noisy, leading to fuzzy segmentations in Fig. 5(c) and making 3D object extraction less robust. In contrast, our method successfully identifies the different objects in the scene in separate object slots.
@@ 4.3 Applications
Since our method can segment the 3D scenes into individual objects, the 3D reconstructions can be used in downstream applications. An object can be extracted by conditioning the rendering of the model in Eq. 1 to only render the associated object slot. This lets a user perform editing operations conditioned on the object slots. As shown in Fig. 7(a), we can remove the bonsai from the reconstructed scene. This operation reveals uncovered geometry under the removed object, which can be fuzzy as it was not supervised during NeRF training. This can easily be enhanced in post-processing by inpainting [30, 40]. Additionally, in Fig. 7(b), we identify objects and change their colors independently. We can also assemble objects from several scenes and render them jointly, as illustrated by Fig. 7(c).
4.4 Additional results on the LERF [18] dataset
We present 3D segmentation results for scenes from the LERF [18] dataset in Fig. 6. We further observe that our method generalizes well to the complex scenes in that dataset. In some cases, high ambiguity in the mask generation can lead to artifacts or fuzzy segmentations, as can be seen at the bottom right of the Waldo kitchen scene.
@@ 5 Conclusion
We have shown that we can successfully learn a meaningful 3D object field from inconsistent 2D masks. We match the channels of our object field with the supervision masks, and regularize the 3D field to produce a spatially consistent 3D segmentation. This is done without requiring semantic labels, which would limit the generality of our approach.
One current limitation is that non contiguous objects may, in rare occasions, be assigned the same label. This is mitigated by our regularization and in practice, such objects can still trivially be separated. Another limitation is that we can currently only handle static scenes. Future work will focus on dynamic NeRFs [34] and enforcing consistency of the segmentations over time. Finally, while our current approach relies on NeRF rendering, it would be interesting to extend it to other representations such as 3D gaussians [17] or physically-based differentiable rendering [16].

### v3

@@ 4.2 Ablation study
To validate the impact of the different components of our method, we perform an ablation study. We remove the matching $\gamma$ and instead match masks and objects by choosing the slot that maximizes the affinity $\alpha$ independently from each other, such that the same object slot can be assigned to several masks. This leads to a sharp drop in our evaluation metrics from $79.24\%$ to $57.04\%$ IoU, as the same object slot often captures several objects in Fig. 5(b).
We also ablate the regularization loss $\mathcal{L}_{TV}$ and compare the produced segmentations in Fig. 5(c). We observe a drop in IoU down to $75.53\%$ that can be attributed to the noisy object field learned without our additional regularization. Interestingly, we also observed in our experiments that the addition of $\mathcal{L}_{TV}$ helped prevent different objects from using the same object slot. We view this as a consequence of our regularization, which prevents the optimization from getting trapped in a local minimum. We further observe that without $\mathcal{L}_{TV}$, confidence remains low and noisy, leading to fuzzy segmentations in Fig. 5(c) and making 3D object extraction less robust. In contrast, our method successfully identifies objects in the scene in separate object slots.
@@ 4.3 Applications
Since our method can segment 3D scenes into individual objects, the 3D reconstructions can be used in downstream applications. An object can be extracted by conditioning the rendering of the model in Eq. 1 to only render the associated slot. As shown in Fig. 7(a), we can remove the bonsai from the scene by removing all its parts. This operation reveals uncovered geometry under the removed object, which can be fuzzy as it was not supervised during NeRF training. This can easily be enhanced in post-processing by inpainting [31, 42]. In Fig. 7(b), we identify objects and change their color channels independently. We can also assemble objects from several scenes and render them jointly, as illustrated by Fig. 7(c).
@@ 5 Conclusion
We have shown that we can successfully learn a meaningful 3D object field from inconsistent 2D masks. We match the channels of our object field with the supervision masks, and regularize the 3D field to produce a spatially consistent 3D segmentation. This is done without requiring semantic labels, which would limit the generality of our approach. One current limitation is that non-contiguous objects may, in rare occasions, be assigned the same label. This is mitigated by our regularization and in practice, such objects can trivially be separated. Although our method is class agnostic, we observe a drop in quality for transparent objects, highly-reflective surfaces, or particularly small objects. Future work will focus on dynamic radiance fields [35] and enforcing consistency of the segmentations over time. Finally, while our current approach relies on NeRFs, it would be interesting to extend it to other representations such as 3DGS [18] or PBDR [17].